# Linear Transformers with Learnable Kernel Functions are Better In-Context Models

### v1

@@ Abstract
**footnotetext: Corresponding author: n.n.balaganskiy@tinkoff.ai
@@ 1 Introduction
Notably, both Linear Transformers Katharopoulos et al. (2020) and SSMs can be described as Recurrent Neural Networks (RNNs) (Chung et al., 2014; Hochreiter and Schmidhuber, 1997), which have their limitations when it comes to managing lengthy dependencies within texts since memory capacity can be overrun as the volume of information increases. Additionally, while the hidden state of RNNs is larger for Linear Transformers than for SSMs, the latter showed higher text modeling quality. The introduction of the Based model (Arora et al., 2023) attempted to address the abovementioned challenges by utilizing a hybrid architecture (Fu et al., 2023a) based on a Linear Transformer with a novel kernel function derived from a Taylor expansion of an exponential function. Arora et al. (2023) demonstrated that the Based model was less prone to performance issues when working with longer content than other models when assessed on the Multi-Query Associative Recall (MQAR) task. Nonetheless, even the Based model experiences a drop in performance when faced with extensive contexts relative to the conventional transformer architecture.
@@
A profound comprehension of the processes occurring within the Based architectures is essential for their advancement. Upon examining how attention scores are distributed, we argue that the kernel function previously adopted in Based cannot be considered optimal, resulting in limitations when dealing with lengthy context and small model capacity. To address this issue, we introduce ReBased (Revisited Based) [^1] Source code is available by the link https://github.com/corl-team/rebased, a novel variation of the Linear Transformer model that improves the use of attention kernels. The crux of our development lies in addressing the inability of Based to disregard specific tokens with zero probability during the attention process. By refining the kernel function and incorporating new architectural modifications, we have created a model that improves accuracy on tasks involving retrieving information from long sequences of tokens while simplifying the calculation of the attention mechanism.
@@ 5.1 Experimental Setup
In language modeling, our second experimental setup leveraged the extensive Pile dataset (Gao et al., 2020) to train a language model (LM). We opted for a sequence length of 4096, a slight increase from the standard value while still ensuring the replication of the architectural framework as presented by Arora et al. (2023) [^2] The experiment details can be found in a blog post and WandB report associated with the main paper.. Note that some hyperparameters such as model dimension and the number of layers were set in order to match the number of model parameters in the initial experiment. Detailed model configuration can be found in Appendix B.
@@ 5.4 Language Modeling
Attention 16.62 3.57 58.57
Based 16.69 3.72 57.21
ReBased 16.45 3.67 56.31
@@
When considering AR perplexity, we observe that there is still a gap between the vanilla Transformer architecture and alternative models, which is aligned with the results on the MQAR dataset. However, we note that ReBased still performed better than Based. Regarding Non-AR perplexity, ReBased outperformed both Based and Transformer architectures, leading to better overall perplexity value.
These results suggest that, despite language modeling perplexity being lower for an alternative to Transformer architectures (Arora et al., 2023; Gu and Dao, 2023), this may be achieved due to better short-term dependency modeling, which does not require learning associative operations necessary to perform In-Context Learning (Olsson et al., 2022). The vanilla Transformer still performs best in terms of its ability to attend to some token in-context.
@@ 5.5 Analysis
Table 3: IoU with attention matrix and ground truth position to retrieve on the MQAR task for $10000$ examples. Detailed experiment setup can be found in 5.5.
@@ 6 Conclusion and Future Work
In this paper, we present ReBased, a novel architecture for sub-quadratic attention computation. For our model, we analyzed the Base architecture and proposed to develop it even further by using polynomial kernels with learnable parameters and adding normalization before the kernel evaluation. While incorporating layer normalization into model training was attempted previously (Dehghani et al., 2023), our method integrates this normalization directly into the kernel function. With this simple architectural change, we achieved results that outperformed Based on MQAR and language modeling with the Pile dataset tasks.
We analyzed the internal representations of ReBased, Based, and vanilla attention modules, and concluded that ReBased resembles attention more than Based. Notably, while Based uses a Taylor expansion of an exponential function, a ReBased kernel function is different from the exponent but shows better performance. Our research suggests that using a second-order polynomial might be insufficient for the best performance, and indicates that more sophisticated learnable kernels could be utilized to improve the performance of trained models. Normalization could further improve various kernel functions. This highlights a need for researchers to revisit classical kernel-based methods with the goal of enhancing their adaptability and efficiency.
Our findings reveal a significant disparity in handling the MQAR task between attention-based models and others such as Based, specifically as sequence lengths increase. Attention models excel on longer sequences, significantly outperforming their non-attention counterparts. These results highlight the necessity of further research into strategies that could bridge this gap in order to reach the performance of attention-based methods. Perhaps the superior aspects of attention mechanisms could be matched or exceeded by other models, especially on tasks that require associative recall, such as machine translation Vardasbi et al. (2023). Future research could give insight into this, leading to improved models for processing long sequences on various NLP tasks.
@@ 7 Limitations
Appendix A Details for the MQAR dataset experiments.
In our experiments, we use the code from the official MQAR repository (Arora et al., 2023) [^3] https://github.com/HazyResearch/zoology. However, we modify the attention model from the one reported in Arora et al. (2023), as we found it more stable (see Figure 7). We can see that replacing the first attention layer is beneficial for performance. RWKV performs better when we do not replace the first layer, which is why we use we use two RWKV layers in our main experiment (see Figure 1). We report the number of trainable parameters in Table 4.
@@
Appendix B Pile Dataset Experiment Details
@@
As in Arora et al. (2023), we use more hyperparameters in the Based/ReBased models then in the Attention baseline. The number of layers and head dim reported in Tables 7 and 8. We use a hybrid architecture for the Based/ReBased models where we use short convolution as a mixer for every odd-numbered layer.
@@
Model # Parameters
Attention 126 807 552
Based 156 955 392
ReBased 156 992 256
Table 9: Parameters count for the pile experiment. See Appendix B.
Appendix C Additional Analysis
@@
Figure 10: Analysis of the layer normalization parameters. Mean value of the scale parameter (gamma) tends to the same value of about $0.13$, and the bias parameter (beta) tends to 00. See Section C.

### v2

@@ Abstract
Linear Transformers with Learnable Kernel Functions are Better In-Context Models
[…]
@@ 1 Introduction
Notably, both Linear Transformers Katharopoulos et al. (2020) and SSMs can be described as Recurrent Neural Networks (RNNs) (Chung et al., 2014; Hochreiter and Schmidhuber, 1997), which have their limitations when it comes to managing lengthy dependencies within texts since memory capacity can be overrun as the volume of information increases. Additionally, while the hidden state of RNNs is larger for Linear Transformers than for SSMs, the latter showed higher text modeling quality. The introduction of the Based model (Arora et al., 2024) attempted to address the abovementioned challenges by utilizing a hybrid architecture (Fu et al., 2023a) based on a Linear Transformer with a novel kernel function derived from a Taylor expansion of an exponential function. Arora et al. (2024) demonstrated that the Based model was less prone to performance issues when working with longer content than other models when assessed on the Multi-Query Associative Recall (MQAR) task. Nonetheless, even the Based model experiences a drop in performance when faced with extensive contexts relative to the conventional transformer architecture.
@@
A profound comprehension of the processes occurring within the Based architectures is essential for their advancement. Upon examining how attention scores are distributed, we argue that the kernel function previously adopted in Based cannot be considered optimal, resulting in limitations when dealing with lengthy context and small model capacity. To address this issue, we introduce ReBased (Revisited Based), a novel variation of the Linear Transformer model that improves the use of attention kernels. The crux of our development lies in addressing the inability of Based to disregard specific tokens with zero probability during the attention process. By refining the kernel function and incorporating new architectural modifications, we have created a model that improves accuracy on tasks involving retrieving information from long sequences of tokens while simplifying the calculation of the attention mechanism.
@@ 5.1 Experimental Setup
In language modeling, our second experimental setup leveraged the extensive Pile dataset (Gao et al., 2020) to train a language model (LM). We opted for a sequence length of 4096, a slight increase from the standard value while still ensuring the replication of the architectural framework as presented by Arora et al. (2024) [^1] The experiment details can be found in a blog post and WandB report associated with the main paper.. Note that some hyperparameters such as model dimension and the number of layers were set in order to match the number of model parameters in the initial experiment. Detailed model configuration can be found in Appendix C.
@@ 5.4 Language Modeling
Attention 11.98 3.07 33.95
Based 12.99 3.27 37.02
ReBased 12.90 3.25 36.73
@@
We note that ReBased model performs better than Based on both AR and non-AR tokens leading to lower overall perplexity. This can be considered as a sign of better In-Context Learning performance. In the next section we consider few-shot setup on several tasks to validate
When considering AR perplexity, we observe that there is still a gap between the vanilla Transformer architecture and alternative models, which is aligned with the results on the MQAR dataset. However, we note that ReBased still performed better than Based. Regarding Non-AR perplexity, ReBased outperformed both Based architectures, leading to better overall perplexity value. Note that attention has slightly more trainable parameters, see Appendix C for more details.
These results suggest that, despite language modeling perplexity being lower for an alternative to Transformer architectures (Arora et al., 2024; Gu and Dao, 2023), this may be achieved due to better short-term dependency modeling, which does not require learning associative operations necessary to perform In-Context Learning (Olsson et al., 2022). The vanilla Transformer still performs best in terms of its ability to attend to some token in-context.
5.5 Few-Shot Performance
To further explore the ability of the model to improve results on real-world scenarios we validate trained Based and ReBased models on a common few-shot benchmarks from the LM Evaluation Harness Gao et al. (2023) benchmark and SuperGLUE Wang et al. (2019). Results are presented in Table 3 and Table 4. ReBased outperforms Based on the most of the tasks.
Architecture Winogrande Piqa Hellaswag Arc-E Arc-C Macro
Based 50.4 62.1 30.8 40.4 22.9 41.3
ReBased 54.6 62.8 30.7 41.0 21.5 42.1
Table 3: 1-shot performance on tasks from LM evaluation harness benchmark Gao et al. (2023). See Section 5.5 for more details.
Architecture WSC WIC RTE Record (F1/EM) MultiRC Copa BoolQ
Based 55.8 46.5 47.6 62.7/62.1 51.5 66.0 48.3
ReBased 56.7 46.9 53.1 62.8/62.2 51.9 67.0 52.0
Table 4: 1-shot performance on SuperGLUE benchmark Wang et al. (2019). See Section 5.5 for more details.
@@ 5.6 Analysis
Table 5: IoU with attention matrix and ground truth position to retrieve on the MQAR task for $10000$ examples. Detailed experiment setup can be found in 5.6.
@@ 6 Conclusion and Future Work
In this paper, we present ReBased, a novel architecture for sub-quadratic attention computation. For our model, we analyzed the Base architecture and proposed to develop it even further by using polynomial kernels with learnable parameters and adding normalization before the kernel evaluation. While incorporating layer normalization into model training was attempted previously (Henry et al., 2020), our method integrates this normalization directly into the kernel function. With this simple architectural change, we achieved results that outperformed Based on MQAR and language modeling with the Pile dataset tasks.
We analyzed the internal representations of ReBased, Based, and vanilla attention modules, and concluded that ReBased resembles attention more than Based. Notably, while Based uses a Taylor expansion of an exponential function, a ReBased kernel function is different from the exponent but shows better performance. Our research suggests that using a second-order polynomial might be insufficient for the best performance, and indicates that more sophisticated learnable kernels could be utilized to improve the performance of trained models. Normalization could further improve various kernel functions. This highlights a need for researchers to revisit kernel-based methods with the goal of enhancing their adaptability and efficiency.
Our findings reveal a disparity in handling the MQAR task between attention-based models and others such as Based, specifically as sequence lengths increase. Attention models excel on longer sequences, significantly outperforming their non-attention counterparts. These results highlight the necessity of further research into strategies that could bridge this gap in order to reach the performance of attention-based methods. Perhaps the superior aspects of attention mechanisms could be matched or exceeded by other models, especially on tasks that require associative recall, such as machine translation Vardasbi et al. (2023). Future research could give insight into this, leading to improved models for processing long sequences.
@@ 7 Limitations
Appendix A Details for the MQAR dataset experiments
In our experiments, we use the code from the official MQAR repository (Arora et al., 2024) [^2] https://github.com/HazyResearch/zoology. However, we modify the attention model from the one reported in Arora et al. (2024), as we found it more stable (see Figure 6). We can see that replacing the first attention layer is beneficial for performance. RWKV performs better when we do not replace the first layer, which is why we use we use two RWKV layers in our main experiment (see Figure 1). We report the number of trainable parameters in Table 6.
@@ Appendix A Details for the MQAR dataset experiments
Appendix B Stability
Appendix C Pile Dataset Experiment Details
Model # Parameters
Attention 151 880 448
Based 147 542 016
ReBased 147 548 928
Table 8: Parameters count for the pile experiment. See Appendix C.
@@ Appendix C Pile Dataset Experiment Details
As in Arora et al. (2023), we use more hyperparameters in the Based/ReBased models then in the Attention baseline. The number of layers and head dim reported in Tables 10 and 11. We use a hybrid architecture for the Based/ReBased models where we use short convolution as a mixer for every odd-numbered layer.
@@
Appendix D Additional Analysis
@@
Figure 10: Analysis of the layer normalization parameters. Mean value of the scale parameter (gamma) tends to the same value of about $0.13$, and the bias parameter (beta) tends to 00. See Section D.